OpenStudy (anonymous):

Hi kindly help ln(2x+3)>0

OpenStudy (anonymous):

solve for x :D ?

OpenStudy (anonymous):

lsending ln to right side of equation gives 2x+3>1 then u can directly solve for x hope it helped :)

OpenStudy (anonymous):

Hint ln x=0 e^(ln x)=e^0 x=1